# Real-world Troublemaker: A Novel Track Testing Framework for Automated Driving Systems in Safety-critical Interaction Scenarios

### v1

@@ III-B Maneuver Challenge Improvement of Interactive Critical Scenarios
To characterize participants state transition, a simplified discrete model is used to represent the vehicle’s kinematics in this study.
@@ III-B1 Risk Interaction Utility Function Design
According to (2), it is necessary to design appropriate metrics to enhance the maneuver challenge, thereby increasing the frequency of events of interest. In this subsection, a multi-attribute reward function is proposed, which serves as one of the core components in interaction scenario generation. While ensuring the generation of collision scenarios, the behavior of object targets should also be made more reasonable and realistic. The interaction utility function primarily considers three aspects: hazard, smoothness, and compliance. It is formulated as a linear combination of the costs defined above, expressed as:
@@
Hazard: To maximize the vehicle collision risk, in this study, the maneuver challenge is estimated by post encroachment time (PET), which calculates the time gap between one actor and another actor in a designed collision point. As discussed in [32], PET is one of the most widely used indices of hazard evaluation, and it is defined as
@@
Figure 3: Unprotected left turn scenario interaction risk illustration.
@@
where the indicator function is negative to represent a yielding behavior by vehicle A, and positive to indicate that vehicle A is rushing. When the function falls within the range of the maneuver challenge, an unsafe collision risk arises.
@@
where m and n represent the boundary values for the safe and unsafe regions in the two scenarios, respectively. $\psi$ represents the PET.
@@
Compliance: To ensure that the object targets do not exhibit unreasonable collision behaviors while performing the test tasks, a predefined test route is established. The expectation is that the vehicle adheres to the road traffic rules throughout the test.
@@ III-B2 Dynamic Interaction Game Strategy Design Based on Stackelberg Equilibrium
In a strongly interactive scenario, a dynamic game problem occurs between the VUT and the object target, which is a typical leader-follower game. The VUT is treated as the follower in the leader-follower game, while the object target serves as the leader. Through adversarial gameplay, both entities make optimal conflict strategies, with each influencing the other’s decision-making. This yields:
$\begin{array}[]{l}(\mathbf{u}^{L*},\sigma^{*})=\underset{\mathbf{u}^{L},\sigma}{\min}\underset{\mathbf{u}^{F}\in\gamma^{2}(\mathbf{u}^{L},\sigma)}{\min}J_{L}(\mathbf{x}^{L},\mathbf{x}^{F},\mathbf{u}^{F},\mathbf{u}^{L},\sigma)\\ \gamma^{2}(\mathbf{u}^{L},\sigma)\\ =\left\{\zeta\in\Phi^{2}:J_{F}(\mathbf{x}^{L},\mathbf{x}^{F},\mathbf{u}^{L},\sigma,\zeta)\leq J_{F}(\mathbf{x}^{L},\mathbf{x}^{F},\mathbf{u}^{L},\sigma,\mathbf{u}^{F}),\right.\\ \quad\quad\left.\forall\mathbf{u}^{F}\in\Phi^{2}\right\}\\ \mathrm{s.t.}\quad\sigma\in\left\{\text{rush},\text{yield}\right\},u_{ij}\in[a_{\text{min}},a_{\text{max}}],v_{ij}\in[0,v_{\text{max}}]\end{array}$ (13)
@@
Clearly, the follower’s objective function is a typical quadratic optimization problem, and the constraints are affine within the feasible domain. Therefore, the follower’s optimal control problem (OCP) is a typical convex optimization problem. Henceforth, the bi-level optimization (X) can be reformulated as a single level problem by using Karush-Kuhn-Tucker (KKT) condition:
@@
The optimization objectives for VUT and the object target are based on the designed reward function, where VUT aims to optimize safety, rule adherence, and smoothness, while the target vehicle optimizes adversarial, rule adherence, and smoothness. The problem is solved using the Cadasi solver to obtain the optimal strategy for the target vehicle. Furthermore, during a single test, the VUT and the target vehicle do not engage in an unrestricted adversarial game, especially when both VUT and the target vehicle adopt a wait-and-yield strategy, which could lead to a deadlock situation. To avoid invalid adversarial interactions, we introduce an adversarial rationality check function. If the game duration M exceeds the maximum interaction threshold $\delta$, the game is terminated. The algorithm for generating Adversarial Game Scenarios is summarized in Algorithm 1.
@@ III-B3 Motion Planning and Control
The motion planning and control problem is decoupled into lateral and longitudinal multi-objective optimization problems along the reference path. Based on the upper-level game results, the optimal state sequence for the target vehicle in the future decision time domain is obtained, and the decision result is represented as a sequence of state points in the frenet coordinate system. To obtain the actual trajectory of the target vehicle and complete the tracking control, a sample based optimization method is used to find an optimal trajectory that satisfies the motion constraints for both lateral and longitudinal directions. The optimal driving path sequence and speed sequence are derived separately, and the trajectories are combined to form the actual driving trajectory. Taking the future state points from the game result as the target, a space sampling method is performed between two adjacent state points. SL graph and ST graph are designed to form the lateral and longitudinal drivable corridors, respectively. The surrounding obstacles (VUT, other targets) are projected onto the graphs, and the reference trajectory that satisfies the constraints and avoids collisions is selected based on the evaluation function. After solving the lateral and longitudinal planning problems, the optimal state sequences are obtained, and a safe, comfortable, and stable optimal motion sequence is synthesized in the frenet coordinate system as the final output trajectory. This trajectory is then transmitted to the target vehicle’s control system via the vehicle-cloud link, where longitudinal control is performed using PID, and lateral control is achieved using Stanley for tracking.
III-C Risk Scenario Exposure Enhancement Method Based on Importance Sampling
@@
However, as shown in Fig.4, the exposure of high-risk scenarios is relatively low. Yet, as indicated by (2), these scenarios contribute more significantly to the occurrence of collision events. Therefore, in the modeling process of decision-making, while accurately reproducing real-world active interaction scenarios, we aim to increase the probability of the event of interest occurring. To achieve this, we employ the importance sampling theorem to increase the exposure rate of high-risk scenarios, thereby accelerating the testing process.
@@
where n denotes the total number of tests, $\psi_{i}\sim q(\psi)$ denotes that the testing scenarios are sampled according to the distribution.
@@
The proposed framework is implemented at the Intelligent Connected Vehicle Evaluation Base at Tongji University. The simulated urban block, as shown in Fig. 5, demonstrates the implementation of the Troublemaker system, including the testing monitoring platform, cloud control center, 5G network, test targets, and VBOX physical setup. The testing monitoring platform is located at the command center of the testing area, where test personnel issue remote test instructions. The cloud control platform generates test scenarios, controls the motion of the test targets, and handles scenario resets based on the instructions. The optimal trajectory of the test targets is then sent to the targets for execution. Meanwhile, VBOX (vehicle state data) and test target information are uploaded in real-time to the cloud control center. The monitoring center can access real-time video stream data and test data from the test field. Data transmission is carried out using 5G and fiber-optic networks. The communication middleware between the test targets and the cloud uses the Message Queuing Telemetry Transport (MQTT) protocol for data transmission. The cloud listens to specific topics to receive vehicle state data and, through a third-party algorithm, sends the obtained vehicle trajectories back to the vehicles via MQTT. The 5G communication network serves as a dedicated network for the test field, with 5G base stations deployed within the testing area. The 5G private network operates in the frequency band of 949.5–959.5 MHz. The cloud-controlled target vehicle is equipped with a self-developed cloud-controlled chassis system, which includes a high-precision positioning module, a 5G communication terminal (CPE), and a VCU. The target vehicle transmits real-time position and orientation data at a frequency of 20 Hz to the cloud control platform via the 5G CPE. The vehicle tracks the cloud-provided trajectory and sends control commands via the CAN bus to the chassis for execution.
@@ V-A2 Test Method for Comparison
The predefined baseline test method, which is the typical testing method currently used at the test site, is adopted as the comparison test method. The relevant parameter values are set according to the site conditions and are described as follows. The VUT drives within the lane and makes a left turn through the intersection. When the pre-collision time between the two vehicles reaches the interval of 4.5 to 5.5 seconds, the object target continues to move straight through the intersection at a constant speed of 3 m/s.
@@ V-A3 Evaluation Metrics
  • Hausdorff Distance of Object Target Velocity: The Hausdorff distance between the object targets’ velocity curves across multiple trials has been widely used in time-series trajectory data to measure the similarity and diversity of different sets of time-series data. In this scenario, since the object targets’ behavior strategy is primarily reflected in their driving velocity, this metric is employed to characterize the diversity of interaction behaviors [34], [35]. A larger Hausdorff distance indicates greater diversity in the interaction behaviors generated across multiple trials. The specific calculation method is provided in (21).
@@ V-B Result and Discussion
Fig. 14 shows that we randomly generated 20 sets of test conditions based on both the empirical distribution and the proposal distribution, and conducted real-world testing based on these conditions. The results indicate that troublemaker-II can generate more severe conflicts (0.9), which is 3.5 times higher than the real-world data distribution (0.2). Another interesting phenomenon observed in the tests is that as the occurrence of high-risk scenarios increases, we find that despite the object target adopting different decision strategies (yielding or trying to overtake), due to the small PET with the object target and the close actual distance between the two vehicles, the VUT tends to take a braking and evasion strategy more often, which aligns with common sense.
@@
The interaction process is dynamic and continuous. To realistically showcase the actual interaction performance during the testing process, we have also made some real-world test video clips available online. Refer to the supplementary videos:
@@ VI Conclusion and future research
Future work will focus on refining the proposal function in importance sampling to dynamically adjust the test difficulty based on the evaluation results, achieving iteratively refined tests tailored to the performance of the ADS. Additionally, we plan to extend the framework to support more complex interaction scenarios, including mixed human-vehicle traffic environments, to further validate its robustness and scalability in real-world applications.

### v2

@@ III-B Maneuver Challenge Improvement Through a Dynamic Game Framework
To characterize participants state transition, a simplified discrete model is used to represent the vehicle’s kinematics.
@@ III-B1 Risk Interaction Utility Function Design
A multi-attribute utility function is proposed as a core component in generating interaction scenarios. While ensuring the creation of collision scenarios, it also aims to make the behavior of object targets more rational and realistic. The interaction utility function primarily considers three aspects: hazard, smoothness, and compliance. It is formulated as a linear combination of the costs defined above, expressed as:
@@
Hazard: To maximize vehicle collision risk, maneuver challenge is estimated by post encroachment time (PET), which calculates the time gap between an actor and another actor in a collision point. As discussed in [35], PET is one of the most widely used indices of hazard evaluation, and it is defined as
@@
Figure 3: The interactive risk illustration.
@@
where the indicator function is negative, representing a yielding behavior by vehicle A, and positive, indicating that vehicle A is rushing. When the function value falls within the range of the maneuver challenge, it signifies an unsafe collision risk, as shown in Fig. 4.
@@
where m and n represent the boundary values for the safe and unsafe regions. $\psi$ represents the PET.
@@
Compliance: To prevent object targets from displaying unrealistic collision behaviors during the test tasks, a predefined test route is established. The expectation is that the vehicle will follow traffic rules throughout the test.
@@ III-B2 Dynamic Interaction Game Strategy Design Based on Stackelberg Equilibrium
In a strongly interactive scenario, a dynamic game problem occurs between the VUT and the object target, which is a typical leader-follower game. The VUT is treated as the follower, while the object target serves as the leader. Through adversarial interaction, both entities make optimal conflict strategies, with each influencing the other’s decision-making. This yields:
$\begin{array}[]{l}(\mathbf{u}^{L*},\sigma^{*})=\underset{\mathbf{u}^{L},\sigma}{\arg\min}\underset{\mathbf{u}^{F}\in\gamma^{2}(\mathbf{u}^{L},\sigma)}{\min}J_{L}(\mathbf{x}^{L},\mathbf{x}^{F},\mathbf{u}^{F},\mathbf{u}^{L},\sigma)\\ \gamma^{2}(\mathbf{u}^{L},\sigma)\\ =\left\{\zeta\in\Phi^{2}:J_{F}(\mathbf{x}^{L},\mathbf{x}^{F},\mathbf{u}^{L},\sigma,\zeta)\leq J_{F}(\mathbf{x}^{L},\mathbf{x}^{F},\mathbf{u}^{L},\sigma,\mathbf{u}^{F}),\right.\\ \quad\quad\left.\forall\mathbf{u}^{F}\in\Phi^{2}\right\}\\ \mathrm{s.t.}\quad\sigma\in\left\{\text{rush},\text{yield}\right\},u_{ij}\in[a_{\text{min}},a_{\text{max}}],v_{ij}\in[0,v_{\text{max}}]\end{array}$ (13)
@@
Clearly, the follower’s objective function is a typical quadratic optimization problem, and the constraints are affine within the feasible domain. Therefore, the follower’s optimal control problem (OCP) is a typical convex optimization problem. Henceforth, the bi-level optimization can be reformulated as a single level problem by using Karush-Kuhn-Tucker (KKT) condition:
@@
The optimization objectives for VUT and the object target are based on the designed reward function, where VUT aims to optimize safety, rule adherence, and smoothness, while the object target optimizes adversarial, rule adherence, and smoothness. The problem is solved using the Cadasi solver to obtain the optimal strategy for the target vehicle. Furthermore, during a single test, the VUT and the target vehicle do not engage in an unrestricted adversarial game, especially when both VUT and the target vehicle adopt a wait-and-yield strategy, which could lead to a deadlock situation. To avoid invalid adversarial interactions, we introduce an adversarial rationality check function. If the game duration M exceeds the maximum interaction threshold $\delta$, the game is terminated. The algorithm for generating Adversarial Game Scenarios is summarized in Algorithm 1.
@@ III-B3 Motion Planning and Control
The motion planning and control problem is decoupled into lateral and longitudinal OCP along the reference path. Based on the upper-level game results, the optimal state sequence for the target vehicle in the future decision time domain is obtained, and the decision result is represented as a sequence of state points in the frenet coordinate system. To obtain the actual trajectory of the target vehicle and complete the tracking control, a sample based optimization method is used to find an optimal trajectory that satisfies the motion constraints for both lateral and longitudinal directions. The optimal driving path sequence and speed sequence are derived separately, and the trajectories are combined to form the actual driving trajectory. Taking the future state points from the game result as the target, a space sampling method is performed between two adjacent state points. SL graph and ST graph are designed to form the lateral and longitudinal drivable corridors, respectively. The surrounding obstacles (VUT, other targets) are projected onto the graphs, and the reference trajectory that satisfies the constraints and avoids collisions is selected based on the evaluation function. After solving the lateral and longitudinal planning problems, the optimal state sequences are obtained, and a safe, comfortable, and stable optimal motion sequence is synthesized in the frenet coordinate system as the final output trajectory. This trajectory is then transmitted to the target vehicle’s control system via the vehicle-cloud link, where longitudinal control is performed using PID, and lateral control is achieved using Stanley for tracking.
III-C Risk Scenario Exposure Enhancement Based on Importance Sampling
@@
The exposure frequency of safety-critical scenarios is relatively low, but these scenarios contribute more significantly to the occurrence of collision events. Therefore, we aim to increase the probability of the event of interest occurring. To achieve this, we employ the importance sampling theorem to increase the exposure rate of high-risk scenarios, thereby accelerating the testing process.
@@
where n denotes the total number of tests, $\psi_{i}\sim q(\psi)$ denotes the sampled scenarios according to the distribution.
@@
The proposed framework has been implemented at the Intelligent Connected Vehicle Evaluation Base of Tongji University. As illustrated in Fig. 5, the simulated urban block demonstrates the comprehensive implementation of the Real-world Troublemaker system, which integrates multiple components: a testing supervision platform, cloud control centre, 5G network infrastructure, object targets, and VBOX. The testing supervision platform, strategically located at the command center of the testing area, serves as the central hub where test personnel initiate and monitor remote test instructions. The cloud control platform plays a pivotal role in generating dynamic test scenarios, orchestrating the motion of object targets, and managing scenario resets in response to received instructions. Following computational processing, the optimal trajectory for object targets is transmitted for execution. Concurrently, critical data streams including VBOX (vehicle state data) and object target information are continuously uploaded to the cloud control center in real-time. The testing supervision platform maintains comprehensive access to both real-time video streams and test data from the field, facilitated by a robust 5G and fiber-optic network infrastructure. The communication architecture employs the message queuing telemetry transport (MQTT) protocol as the middleware for data transmission between targets and the cloud. The cloud system actively monitors specific topics to acquire vehicle state data, which is then processed through third-party algorithms to generate vehicle trajectories. These trajectories are subsequently transmitted back to the vehicles via the MQTT protocol. The dedicated 5G communication network, operating within the 949.5–959.5 MHz frequency band, has been specifically deployed for the test field, with 5G base stations strategically positioned throughout the testing area. The cloud-controlled target vehicle is equipped with an advanced, self-developed cloud-controlled chassis system, incorporating three core components: a high-precision positioning module, a 5G customer premises equipment (CPE), and a sophisticated vehicle control unit. The object targets maintain continuous communication with the cloud center, transmitting real-time position and orientation data at a frequency of 20 Hz through the 5G CPE. These targets precisely follow the cloud-generated trajectories, with control commands being executed via the CAN bus interface to the vehicle chassis.
@@ V-A2 Test Method for Comparison
The predefined baseline test method, which is the typical testing approach currently used at the test site, serves as the comparison method. Relevant parameter values are set based on the site conditions and are described as follows. The VUT drives within its lane and makes a left turn through the intersection. When the pre-collision time between the two vehicles reaches an interval of 4.5 to 5.5 seconds, the object target continues to move straight through the intersection at a constant speed of 3 m/s.
@@ V-A3 Evaluation Metrics
  • Hausdorff Distance of Object Target Velocity: The Hausdorff distance between the object targets’ velocity curves across multiple trials has been widely used in time-series trajectory data to measure the similarity and diversity of different sets of time-series data. In this scenario, since the object targets’ behavior strategy is primarily reflected in their driving velocity, this metric is employed to characterize the diversity of interaction behaviors [37]. A larger Hausdorff distance indicates greater diversity in the interaction behaviors generated across multiple trials. The specific calculation is provided in (21).
@@ V-B Result and Discussion
Fig. 14 shows that we randomly generated 20 sets of test conditions based on both the empirical distribution and the proposal distribution, and conducted real-world testing based on these conditions. The results indicate that troublemaker can generate more severe conflicts (0.9), which is 3.5 times higher than the real-world data distribution (0.2). Another interesting phenomenon observed in the tests is that as the occurrence of high-risk scenarios increases, we find that despite the object target adopting different decision strategies (yielding or trying to overtake), due to the small PET with the object target and the close actual distance between the two vehicles, the VUT tends to take a braking and evasion strategy more often, which aligns with common sense.
@@
The interaction process is dynamic and continuous. To realistically showcase the testing process, we have also made some real-world test video clips available online. Refer to the supplementary videos:
@@ VI Conclusion and future research
Future work will refine the proposal function in importance sampling to dynamically adjust test difficulty, enabling iterative tests tailored to ADS performance. We also plan to extend the framework to handle more complex scenarios, including mixed human-vehicle traffic, to further validate its robustness and scalability in real-world applications.